# EventFlow: Forecasting Temporal Point Processes with Flow Matching

### v2

@@ 15       $\gamma_{s_{k}}\leftarrow\gamma_{s_{k-1}}+h_{k}v_{\theta}(\gamma_{s_{k-1}},s_{k-1},e_{\mathcal{H}})$
Once $v_{\theta}$ is learned, we may sample from the model by drawing a reference sequence $\gamma_{0}\sim\mu_{0}$ and solving the corresponding ODE parametrized by $v_{\theta}$. More concretely, we first fix a number of events $n$. When seeking to unconditionally generate new sequences from the underlying data TPP $\mu_{1}$, we simply sample $n$ from the empirical event count distribution $\mu_{1}(n)$. For conditional tasks, we draw $n\sim p_{\phi}(n\mid\mathcal{H})$ from the learned conditional distribution over event counts. Next, we draw $n$ initial events, corresponding to $s=0$, by sampling and sorting $t_{0}^{1},\dots,t_{0}^{n}\sim q$. In practice, we use $q=\mathcal{N}(0,I_{n})$ as we normalize our sequences into the range $[-1,1]$ during training and sampling (followed by renormalization to the data scale). Since we have fixed $n$, we may view this initial draw as a vector $\gamma_{0}=[t_{0}^{1},\dots,t_{0}^{n}]\in\mathcal{T}^{n}$. This event sequence $\gamma_{0}$ then serves as the initial condition for the system of ODEs $d\gamma_{s}=v_{\theta}(\gamma_{s},s)\,\mathrm{d}s$ which can be solved numerically. In our experiments, we use the forward Euler scheme, i.e., we specify a discretization $\{0=s_{0}<s_{1}<\dots<s_{K}=1\}$ of the flow time (in practice, uniform) and recursively compute
$\gamma_{s_{k}}=\gamma_{s_{k-1}}+h_{k}v_{\theta}(\gamma_{s_{k-1}},s_{k-1})\qquad k=1,2,\dots,K$ (8)
where $h_{k}=s_{k}-s_{k-1}$ is a step size. While other choices are possible, we found that this simple scheme was sufficient as the model sample paths are typically close to linear. See Algorithm 2.
@@ 5 Experiments
We study our proposed EventFlow model under two settings. The first is a conditional task, where we seek to forecast both the number of future events and their times of occurrence, over a particular horizon, given a history. The second is an unconditional task, where we aim to learn a representation of the underlying TPP distribution from empirical observations and generate new sequences from this distribution. This second task can be viewed as a special case of the first with no observed history. We evaluate our model across a diverse set of datasets encompassing a wide range of possible point process behaviors. We evaluate our model on seven real-world datasets, which are a standard benchmark for modeling unmarked TPPs (Shchur et al., 2020b; Bosser & Taieb, 2023; Lüdke et al., 2023). We also evaluate on a collection of six synthetic datasets (Omi et al., 2019). See Appendix A.
@@
We selected a set of baselines consisting of a set of diverse and highly performant neural TPP models. For an intensity-based method, we compare against the Neural Hawkes Process (NHP) (Mei & Eisner, 2017), a well-known and very widely-used neural TPP model. We additionally compare against two intensity-free methods, namely the flow-based IFTPP model (Shchur et al., 2020a) and the diffusion-based model of Lin et al. (2022). Lastly, our strongest baseline is the recently proposed Add-and-Thin model of Lüdke et al. (2023), which can be loosely viewed as a non-autoregressive diffusion model. These models use an RNN-based history encoder, with the exception of Add-and-Thin which uses a CNN-based encoder. See Appendix D for additional details.
@@
Evaluating generative TPP models is challenging, as one must take into account both the variable locations and numbers of events. This is particularly challenging for the unconditional setting, where unlike forecasting, we do not have a ground-truth sequence to compare against. Our starting point is a metric (Xiao et al., 2017a) on the space of sequences $\Gamma$, allowing us to measure the distance between two sequences $\gamma=\sum_{k=1}^{n}\delta[t^{k}_{\gamma}]$ and $\eta=\sum_{k=1}^{m}\delta[t^{k}_{\eta}]$ with possibly different numbers of events. Without loss of generality, we assume $n\leq m$, where the distance is given by
@@
where we recall that sequences are supported on $\mathcal{T}=[0,T]$. This distance can be understood either as an $L^{1}$ distance between the counting processes of $\gamma,\eta$ or as a generalization of the $1$-Wasserstein distance to measures of unequal mass. For our unconditional experiment, we require a metric which will capture the distance between the TPP distributions themselves. To do so we use the distance in Equation (9) to calculate an MMD (Gretton et al., 2012; Shchur et al., 2020b). We use an exponential kernel $k(\gamma,\eta)=\exp\left(-d(\gamma,\eta)/(2\sigma^{2})\right)$ with $\sigma$ chosen to be the median distance between all sequences (Shchur et al., 2020b; Lüdke et al., 2023).
@@ 5.1 Forecasting Event Sequences
We first evaluate our model on a multi-step forecasting task. We set a horizon $\Delta T$ for each of our real-world datasets and seek to generate event sequences in the range $[T_{0},T_{0}+\Delta T]$ for some given $T_{0}$, conditioned on the history $\mathcal{H}_{T_{0}}$. Up to a shift, this means we are taking $\mathcal{T}=[0,\Delta T]$. At training time, we uniformly sample $T_{0}\in[\Delta T,T-\Delta T]$. At test time, we sample 50 values of $T_{0}$ for each test set sequence. We then forecast the sequence in $[T_{0},T_{0}+\Delta T]$ and compute the distance (9) between the ground-truth and generated sequences. We note that the distance in Equation (9) is computed using $T_{0}+\Delta T$ rather than $T$ as the maximum event time, as using $T$ would result in a distance which is sensitive to the location of the forecasting window. We further normalize Equation (9) by $\Delta T$. This overall approach is similar to that of Lüdke et al. (2023). For all datasets, the sequences are split into train-validation-test sets, and all metrics are reported on the held-out test set.
We report the results of these experiments in Figure 2. Our proposed EventFlow method obtains the lowest average forecasting error across all datasets, achieving an error which is 20%-53% lower than the nearest baseline. Given that the non-autoregressive models (EventFlow, Add-and-Thin) consistently outperform the autoregressive baselines, this is clear evidence that autoregressive models can struggle on multi-step predictions. This is especially true on the Reddit-C and Reddit-S datasets, which exhibit long sequence lengths.
Table 1: Mean sequence distance per dataset for EventFlow as we vary the NFEs used to simulate the ODE.
NFEs PUBG Red.C Red.S Taxi Twitter Yelp-A Yelp-M
25 $2.0$ $15.8$ $16.0$ 3.23.2{3.2}3.2 1.41.4{1.4}1.4 1.31.3{1.3}1.3 1.91.9{1.9}1.9
@@
In Table 1, we additionally vary the number of function evaluations (NFEs) used at test time, i.e., the number of steps $K$ used in the ODE solver in Algorithm 2. We find that reducing $K$ results in minimal (or no) loss in performance. Even with only a single NFE, EventFlow is able to obtain state-of-the-art performance. This is because, although the vector fields (6) in EventFlow are generally non-linear, we find that our construction produces paths which are approximately so. In contrast, Add-and-Thin uses 100 NFEs, while the NFEs for the autoregressive models scale linearly with the number of sampled events.
@@
PUBG Red.C Red.S Taxi Twitter Yelp-A Yelp-M
@@
Add-and-Thin $\underline{0.43}$ $0.99$ $\underline{0.38}$ $\underline{0.33}$ $\underline{0.60}$ $\mathbf{0.42}$ $\mathbf{0.46}$
EventFlow (ours) $\mathbf{0.40}$ $\mathbf{0.70}$ $\mathbf{0.16}$ $\mathbf{0.28}$ $\mathbf{0.46}$ $\underline{0.56}$ $\underline{0.50}$
To evaluate the predictions of event counts, ${p_{\phi}(n\mid\mathcal{H})}$, we report the mean absolute relative error (MARE) between the true and predicted event counts. This is reported in Table 2. Our model achieves strong performance on this metric as it decouples the event count prediction from the event time prediction, allowing us to train explicitly for each task. The baselines, on the other hand, only model $n$ implicitly. See Appendix E for details and standard deviations.
5.2 Unconditional Generation of Event Sequences
Finally, we evaluate our model on an unconditional generation task, where we aim to generate new sequences from the underlying data distribution. In Table 3 we report MMD values for each of the synthetic and real-world datasets. See Appendix E for results with standard deviations. EventFlow (mean rank: $\mathbf{1.8}$) exhibits uniformly strong performance, obtaining either the best or second best MMD on 11 of the 13 datasets. While IFTPP (mean rank: $\mathbf{2.1}$) performs well on the synthetic datasets, it struggles with the more complex real-world datasets, reflecting its high bias. The Add-and-Thin method (mean rank: $\mathbf{2.4}$) is often similarly strong, but struggles on the SC dataset. While the NHP (mean rank: $\mathbf{3.7}$) can obtain good fits, this appears to be dataset dependent, with weak results on the NSR, SC, and Reddit-S datasets. The diffusion baseline (mean rank: $\mathbf{4.8}$) is our weakest baseline, which is perhaps unsurprising as this model can only be trained to maximize the likelihood of a subsequent event and not the overall sequence likelihood.
@@ 6 Conclusion
In this work, we propose EventFlow, a non-autoregressive generative model for temporal point processes. We demonstrate that EventFlow is able to achieve state-of-the-art results on a multi-step forecasting task and strong performance on unconditional generation. Our work lays a foundation for future flow-based TPPs, and exploring applications in tasks like imputation, marked TPPs, and querying (Boyd et al., 2023) are exciting directions.
@@
In this work, we have focused on modeling the event times, and thus do not experiment with marked TPPs or spatiotemporal TPPs. Extending EventFlow to these tasks is a promising directions. We also note that we do not explicitly enforce that the support of our model is $[0,T]$, which would necessitate moving beyond the Gaussian setting. Moreover, our baseline methods use the neural architectures proposed in the original works, which vary across the methods. Carefully studying the effects of these architecture choices is an important direction.
Acknowledgments
This research was supported by the National Institutes of Health under award R01-LM013344 and by the Hasso Plattner Institute (HPI) Research Center in Machine Learning and Data Science at the University of California, Irvine.
@@ Appendix A Datasets
Our synthetic datasets are adopted from those proposed by Omi et al. (2019). Each of these datasets consists of $1,000$ sequences supported on $\mathcal{T}=[0,100]$. These synthetic datasets are chosen as they exhibit a wide range of behavior, ranging from i.i.d. inter-arrival times to self-correcting processes which discourage rapid bursts of events. We refer to Section 4 of Omi et al. (2019) for details.
@@
We use the set of real-world datasets proposed in Shchur et al. (2020b), which constitute a set of standard benchmark datasets for unmarked TPPs. We refer to Appendix D of Shchur et al. (2020b) for additional details. With the exception of PUBG, these datasets are supported on $\mathcal{T}=[0,24]$, i.e. each sequence corresponds to a single day. For the PUBG dataset, $\mathcal{T}=[0,38]$ corresponds to the maximum length (in minutes) of an online game of PUBG. We note that PUBG has the largest number of sequences (which can lead to slow training), and the Reddit-C and Reddit-S datasets have long sequences (which can lead to slow training and high memory costs).
For selecting the forecasting horizon $\Delta T$, we follow the choice made in Lüdke et al. (2023) to facilitate comparison and reproducibility. We note the initial times $T_{0}\in[\Delta T,T-\Delta T]$ are selected to be at least $\Delta T$ to avoid forecasting sequences with no observations in the history.
@@
Hawkes1 1000 95.4 45.8 [14, 300] [0, 100] {-}-
Hawkes2 1000 97.2 49.1 [18, 355] [0, 100] {-}-
Nonstationary Poisson 1000 100.3 9.8 [71, 134] [0, 100] {-}-
Nonstationary Renewal 1000 98 2.9 [86, 100] [0, 100] {-}-
Stationary Renewal 1000 109.2 38.1 [1, 219] [0, 100] {-}-
Self-Correcting 1000 100.3 0.74 [98, 102] [0, 100] {-}-
PUBG 3001 76.5 8.8 [26, 97] [0, 38] $5$
Reddit-C 1356 295.7 317.9 [1, 2137] [0, 24] $4$
Reddit-S 1094 1129 359.5 [363, 2658] [0, 24] $4$
Taxi 182 98.4 20 [12, 140] [0, 24] $4$
Twitter 2019 14.9 14 [1, 169] [0, 24] $4$
Yelp-Airport 319 30.5 7.5 [9, 55] [0, 24] $4$
Yelp-Miss. 319 55.2 15.9 [3, 107] [0, 24] $4$

### v3

@@ 4 EventFlow
Once $v_{\theta}$ is learned, we may sample from the model by drawing a reference sequence $\gamma_{0}\sim\mu_{0}$ and solving the corresponding ODE parametrized by $v_{\theta}$. We fix a number of events $n$, which can be sampled from the empirical event count distribution $\mu_{1}(n)$ for unconditional generation. For conditional tasks, we draw $n\sim p_{\phi}(n\mid\mathcal{H})$ from the learned conditional distribution over event counts. Next, we draw $n$ initial events, corresponding to $s=0$, by sampling and sorting $t_{0}^{1},\dots,t_{0}^{n}\sim q$. In practice, we use $q=\mathcal{N}(0,I_{n})$ as we normalize our sequences into the range [1,1][-1,1] during training and sampling (followed by renormalization to the data scale). Since we have fixed $n$, we may view this initial draw as a vector $\gamma_{0}=[t_{0}^{1},\dots,t_{0}^{n}]\in\mathcal{T}^{n}$. This event sequence $\gamma_{0}$ then serves as the initial condition for the system of ODEs $d\gamma_{s}=v_{\theta}(\gamma_{s},s)\,\mathrm{d}s$ which can be solved numerically. In our experiments, we use the forward Euler scheme, i.e., we specify a discretization $\{0=s_{0}<s_{1}<\dots<s_{K}=1\}$ of the flow time (in practice, uniform) and recursively compute
$\gamma_{s_{k}}=\gamma_{s_{k-1}}+h_{k}v_{\theta}(\gamma_{s_{k-1}},s_{k-1}),\quad k=1,\dots,K$ (11)
where $h_{k}=s_{k}-s_{k-1}$ is a step size. While other choices are possible, we found that this simple scheme was sufficient as the model sample paths are, qualitatively, typically close to linear. See Algorithm 2.
@@ 5 Experiments
We study our proposed EventFlow model under two settings [^2] Code for our experiments is available at this URL: https://github.com/GavinKerrigan/eventflow. The first is a conditional task, where we seek to forecast both the number of future events and their times of occurrence, over a particular horizon, given a history. The second is an unconditional task, where we aim to learn a representation of the underlying TPP distribution from empirical observations and generate new sequences from this distribution. This second task can be viewed as a special case of the first with no observed history.
We evaluate our model across a diverse set of datasets encompassing a wide range of possible point process behaviors. We evaluate our model on seven real-world datasets, which are a standard benchmark for modeling unmarked TPPs (Shchur et al., 2020b; Bosser and Taieb, 2023; Lüdke et al., 2023). We also evaluate on a collection of six synthetic datasets (Omi et al., 2019). See Appendix A for additional information regarding our datasets.
@@
We selected a set of baselines consisting of a set of diverse and highly performant neural TPP models. For an intensity-based method, we compare against the Neural Hawkes Process (NHP) (Mei and Eisner, 2017), a well-known and very widely-used neural TPP model. We additionally compare against two intensity-free methods, namely the flow-based IFTPP model (Shchur et al., 2020a) and the diffusion-based model of Lin et al. (2022). Lastly, our strongest baseline is the Add-and-Thin model of Lüdke et al. (2023), which can be loosely viewed as a non-autoregressive diffusion model. These models use an RNN-based history encoder, with the exception of Add-and-Thin which uses a CNN-based encoder. See Appendix D for additional details on our baselines.
@@
Evaluating generative TPP models is challenging, as one must take into account both the variable locations and numbers of events. This is particularly challenging for the unconditional setting, where unlike forecasting, we do not have a ground-truth sequence.
Our starting point is a metric (Xiao et al., 2017a) on the space of sequences $\Gamma$, allowing us to measure the distance between two sequences $\gamma=\sum_{k=1}^{n}\delta[t^{k}_{\gamma}]$ and $\eta=\sum_{k=1}^{m}\delta[t^{k}_{\eta}]$ with possibly different numbers of events. Without loss of generality, we assume $n\leq m$, in which case the distance is given by
@@
where we recall that sequences are supported on $\mathcal{T}=[0,T]$. This distance can be understood either as an $L^{1}$ distance between the counting processes of $\gamma,\eta$ or as a generalization of the $1$-Wasserstein distance to measures of unequal mass. For our unconditional experiment, we require a metric that will capture the distance between the TPP distributions themselves. To do so we use the distance in Equation (12) to calculate an MMD (Gretton et al., 2012; Shchur et al., 2020b). We use an exponential kernel $k(\gamma,\eta)=\exp\left(-d(\gamma,\eta)/(2\sigma^{2})\right)$ with $\sigma$ chosen to be the median distance between all sequences in a dataset (Shchur et al., 2020b; Lüdke et al., 2023).
@@ 5.1 Forecasting Event Sequences
We first evaluate our model on a multi-step forecasting task. We set a horizon $\Delta T$ for each of our real-world datasets and seek to generate event sequences in the range $T_{0},T_{0}+\Delta T]$ for some given $T_{0}$, conditioned on the history $\mathcal{H}_{T_{0}}$. Up to a shift, this means we are taking $\mathcal{T}=[0,\Delta T]$. At training time, we uniformly sample $T_{0}\in[\Delta T,T-\Delta T]$. At test time, we sample 50 values of $T_{0}$ for each test set sequence. We then generate one forecast for the sequence in $T_{0},T_{0}+\Delta T]$ and compute the distance (12) between the ground-truth and generated sequences. We note that the distance in Equation (12) is computed using $T_{0}+\Delta T$ rather than $T$ as the maximum event time, as using $T$ would result in a distance which is sensitive to the location of the forecasting window. We further normalize Equation (12) by $\Delta T$. This overall approach is similar to that of Lüdke et al. (2023). For all datasets, the sequences are split into train-validation-test sets, and all metrics are reported on the held-out test set.
We report the results of these experiments in Figure 2. Our proposed EventFlow method obtains the lowest average forecasting error across all datasets, achieving an error which is 20%-53% lower than the nearest baseline. Given that the non-autoregressive models (EventFlow, Add-and-Thin) typically outperform the autoregressive baselines, we see clear evidence that autoregressive models can struggle on multi-step tasks. This is especially true on Reddit-C and Reddit-S which exhibit long sequence lengths.
Table 1: Mean sequence distance for EventFlow as we vary the NFEs used to simulate the ODE.
NFEs PUBG Red.C Red.S Taxi Tw. Yelp-A Yelp-M
25 $2.0$ $15.8$ $16.0$ 3.2{3.2} 1.4{1.4} 1.3{1.3} 1.9{1.9}
@@
PUBG Red.C Red.S Taxi Tw. Yelp-A Yelp-M
@@
Add-Thin $\underline{0.43}$ $0.99$ $\underline{0.38}$ $\underline{0.33}$ $\underline{0.60}$ $\mathbf{0.42}$ $\mathbf{0.46}$
EventFlow $\mathbf{0.40}$ $\mathbf{0.70}$ $\mathbf{0.16}$ $\mathbf{0.28}$ $\mathbf{0.46}$ $\underline{0.56}$ $\underline{0.50}$
In Table 1, we additionally vary the number of function evaluations (NFEs) used at test time, i.e., the number of steps $K$ used in the ODE solver in Algorithm 2. We find that reducing $K$ results in minimal (or no) loss in performance. Even with only a single NFE, EventFlow is able to obtain state-of-the-art performance. This is due to the fact that, although the vector fields (Equation (9)) in EventFlow are generally non-linear, we find qualitatively that our construction produces paths which are approximately so. In contrast, Add-and-Thin uses 100 NFEs, while the NFEs for the autoregressive models scale linearly with the number of generated events.
@@
To evaluate the event count predictions, $p_{\phi}(n\mid\mathcal{H})$}, we report the mean absolute relative error (MARE) between the true and predicted counts in Table 2. Our model achieves strong performance on this metric as it decouples the event count prediction from the event time prediction, training explicitly for each task. The baselines, on the other hand, only model $n$ implicitly. See Appendix E for details and standard deviations.
5.2 Unconditional Generation of Event Sequences
Finally, we evaluate our model on an unconditional generation task, where we aim to generate new sequences from the underlying data distribution without conditioning on an observed history. In Table 3 we report MMD values for each of the synthetic and real-world datasets. MMDs are calculated by sampling 1,000 sequences from each trained model. The MMD values in Table 3 correspond to the mean test set MMD ($\pm$ one standard deviation) across five random seeds. See Appendix E for results with standard deviations.
EventFlow (mean rank: $\mathbf{1.8}$) exhibits uniformly strong performance, obtaining either the best or second best MMD on 11 of the 13 datasets. This is particularly pronounced on the real-world datasets, where we obtain the lowest MMD on 5 of the 7 datasets.
Notably, IFTPP (mean rank: $\mathbf{2.1}$) is one of the strongest methods on the synthetic data, but is relatively weak on the real-world data. We believe that this can be understood in terms of model flexibility. IFTPP uses a simple parametric form (a mixture of log-normals) to parametrize the subsequent event time distribution (see Section 3). This provides a strong inductive bias and allows it to fit simple synthetic distributions very effectively, but limits its capability to model more realistic data sets. EventFlow, in contrast, makes no parametric assumptions, offering greater flexibility. As a result, it may not outperform simple parametric models on very simple synthetic datasets, but this flexibility is precisely what enables strong performance on complex real-world data, which we consider the more meaningful benchmark.
The Add-and-Thin method (mean rank: $\mathbf{2.4}$) is often similarly strong, but struggles on the SC dataset. While the NHP (mean rank: $\mathbf{3.7}$) can obtain good fits, this appears to be dataset dependent, with weak results on the NSR, SC, and Reddit-S datasets. The diffusion baseline (mean rank: $\mathbf{4.8}$) is our weakest baseline, which is perhaps unsurprising as this model can only be trained to maximize the likelihood of a subsequent event and not the overall sequence likelihood.
@@ 6 Conclusion
We introduced EventFlow, a non-autoregressive generative model for temporal point processes. Our approach extends flow matching to the TPP setting by constructing a continuous transformation between a simple reference distribution and the data distribution over event sequences. Concretely, EventFlow generates event sequences by first drawing a sequence of event times from a reference process, and then transporting these events along a learned vector field. Beyond modeling event times, our framework also captures the distribution over the number of events, which we learn via a regularized cross-entropy objective.
Empirically, we demonstrate that EventFlow achieves state-of-the-art performance on multi-step forecasting tasks and competitive results on unconditional generation across a range of standard benchmarks. These results highlight the promise of non-autoregressive, flow-based approaches as a flexible alternative to classical intensity-based models for TPPs.
@@
Our work focuses on modeling event times, and we do not consider marked or spatiotemporal point processes. Extending EventFlow to incorporate marks or spatial structure is a natural and promising direction for future work, though it may require new coupling and interpolation strategies.
Additionally, our current construction does not explicitly enforce support constraints on $[0,T]$, as it relies on Gaussian noising in Equation (6). Addressing this limitation, e.g., by designing constrained flows, could potentially lead to further gains in performance. Finally, our experimental comparisons adopt the neural architectures used in prior work, which differ across baseline methods. A more controlled study of architectural choices, and their interaction with flow-based objectives, would help isolate the gains attributable to the modeling framework itself.
More broadly, we believe our work opens the door to a new class of generative models for point processes based on flow-based ideas, and we hope it encourages further exploration of non-autoregressive approaches.
Acknowledgements
We thank the reviewers for their feedback on improving the paper. This work was supported by National Science Foundation under awards NSF 2505006 and NSF 2425932, by the National Institutes of Health under awards R01-LM013344 and R01CA297869, by the UK EPSRC grant EP/Y037200/1, by the Hasso Plattner Institute (HPI) Research Center in Machine Learning and Data Science at UCI, and by funding support from Google and from SAP.
Supplementary Materials
@@ Appendix A Datasets
Our synthetic datasets are adopted from those proposed by Omi et al. [2019]. Each of these datasets consists of $1,000$ sequences supported on $\mathcal{T}=[0,100]$. These synthetic datasets are chosen as they exhibit a wide range of behavior, ranging from i.i.d. inter-arrival times to self-correcting processes which discourage rapid bursts of events. We refer to Section 4 of Omi et al. [2019] for details.
@@
We use the set of real-world datasets proposed in Shchur et al. [2020b], which constitute a set of standard benchmark datasets for unmarked TPPs. We refer to Appendix D of Shchur et al. [2020b] for additional details. With the exception of PUBG, these datasets are supported on $\mathcal{T}=[0,24]$, i.e. each sequence corresponds to a single day. For the PUBG dataset, $\mathcal{T}=[0,38]$ corresponds to the maximum length (in minutes) of an online game of PUBG. We note that PUBG has the largest number of sequences (which can lead to slow training), and the Reddit-C and Reddit-S datasets have long sequences (which can lead to slow training and high memory costs).
For selecting the forecasting horizon $\Delta T$, we follow the choice made in Lüdke et al. [2023] to facilitate comparison and reproducibility. We note the initial times $T_{0}\in[\Delta T,T-\Delta T]$ are selected to be at least $\Delta T$ to avoid forecasting sequences with no observations in the history.
@@
Hawkes1 1000 95.4 45.8 [14, 300] [0, 100] {-}
Hawkes2 1000 97.2 49.1 [18, 355] [0, 100] {-}
Nonstationary Poisson 1000 100.3 9.8 [71, 134] [0, 100] {-}
Nonstationary Renewal 1000 98 2.9 [86, 100] [0, 100] {-}
Stationary Renewal 1000 109.2 38.1 [1, 219] [0, 100] {-}
Self-Correcting 1000 100.3 0.74 [98, 102] [0, 100] {-}
PUBG 3001 76.5 8.8 [26, 97] [0, 38] 55
Reddit-C 1356 295.7 317.9 [1, 2137] [0, 24] 44
Reddit-S 1094 1129 359.5 [363, 2658] [0, 24] 44
Taxi 182 98.4 20 [12, 140] [0, 24] 44
Twitter 2019 14.9 14 [1, 169] [0, 24] 44
Yelp-Airport 319 30.5 7.5 [9, 55] [0, 24] 44
Yelp-Miss. 319 55.2 15.9 [3, 107] [0, 24] 44